A company produces steel rods. The lengths of the steel rods are normally distributed with a mean of 158.2-cm and a standard deviation of 0.8-cm. For shipment, 9 steel rods are bundled together.

Find the probability that the average length of a randomly selected bundle of steel rods is greater than 158.7-cm.
P(M > 158.7-cm) =

Enter your answer as a number accurate to 4 decimal places. Answers obtained using exact z-scores or z-scores rounded to 3 decimal places are accepted
Answer #1

Solution :

Given that ,

mean = = 158.2

standard deviation = =0.8

P(M >158.7 ) = 1 - p( M< 158.7 )

=1- p P[(x - ) / < (158.7- 158.2) /0.8 ]

=1- P(z < 0.625 )

= 1 - 0.7357 = 0.2643

PROBABILITY = 0.2643